ONE KILOGRAM OF A DIATOMIC GAS IS AT A PRESSURE OF 8 N/m2. THE DENSITY OF THE GAS IS 4 kg/m3 .WHAT IS THE ENERGY DUE TO THERMAL MOTION ?

Grade:12

1 Answers

Arun
5 years ago
Dear Ragini

E = (5/2) nRT
= (5/2) PV
= (5/2) P (m/P)
= (5 x 8 x 104 x 1)/(2 x 4)
= 5 x 104 J
Hope it helps

Regards
Arun (askIITians forum expert)